Use the following information for the next four (4) questions Suppose you work for the Las Vegas Chamber of Commerce, and you want to with 95% confidence, the difference between the percentage of all females who have gone to see an Elvis impersonator and the percentage of all impersonator. A random sample of 100 females includes 53 females who have seen an Elvis impersonator. A random sample of 110 males includes 37 males who have seen an Elvis impersonator estimate males who have gone to see an Elvis 41. For a 95% confidence interval, the z value is (a) 1.96 (b) 1.645 (c) 0.92 (d) 0.65 42. The difference between the sample proportions (females - males) is (a) 1.086 (b) 0.194 (c) 0.211 (d) 0.332
media%2Fe70%2Fe70a9edd-8751-44d1-981c-14
Answer #1

Solution:-

PFemale = 53/100 = 0.53

PMale = 37/110 = 0.3364

41) The zalpha/2 value for the 95% confidence interval is 1.96.

42) b) The difference between the sample proportion is 0.194

Difference = PFemale - PMale

Difference = 0.1936

43) The margin of error is

M.E = z_{\alpha /2}\times \sqrt{\frac{p_{1}(1-p_{1})}{n_{1}}+\frac{p_{2}(1-p_{2})}{n_{2}}}

M.E = 1.96 × 0.06723

M.E = 0.1318

44)

State the hypotheses. The first step is to state the null hypothesis and an alternative hypothesis.

Null hypothesis: P1 = P2
Alternative hypothesis: P1 ≠ P2

Note that these hypotheses constitute a two-tailed test. The null hypothesis will be rejected if the proportion from population 1 is too big or if it is too small.

Formulate an analysis plan. For this analysis, the significance level is 0.05. The test method is a two-proportion z-test.

Analyze sample data. Using sample data, we calculate the pooled sample proportion (p) and the standard error (SE). Using those measures, we compute the z-score test statistic (z).

p = (p1 * n1 + p2 * n2) / (n1 + n2)

p = 0.4286

SE = sqrt{ p * ( 1 - p ) * [ (1/n1) + (1/n2) ] }
SE = 0.06838

z = (p1 - p2) / SE

z = 2.83

where p1 is the sample proportion in sample 1, where p2 is the sample proportion in sample 2, n1 is the size of sample 1, and n2 is the size of sample 2.

Since we have a two-tailed test, the P-value is the probability that the z-score is less than - 2.83 or greater than 2.83

Thus, the P-value = 0.0046

Interpret results. Since the P-value (0.0046) is less than the significance level (0.05), we cannot accept the null hypothesis.

c) H0 is rejected, there is difference in proportions.